# Supervised Graph Contrastive Learning for Gene Regulatory Network

### v3

@@ Abstract
Graph representation learning is effective for obtaining a meaningful latent space utilizing the structure of graph data and is widely applied, including biological networks. In particular, Graph Contrastive Learning (GCL) has emerged as a powerful self-supervised method that relies on applying perturbations to graphs for data augmentation. However, when applying existing GCL methods to biological networks such as Gene Regulatory Networks (GRNs), they overlooked meaningful biologically relevant perturbations, e.g., gene knockdowns. In this study, we introduce SupGCL (Supervised Graph Contrastive Learning), a novel GCL method for GRNs that directly incorporates biological perturbations derived from gene knockdown experiments as the supervision. SupGCL mathematically extends existing GCL methods that utilize non-biological perturbations to probabilistic models that introduce actual biological gene perturbation utilizing gene knockdown data. Using the GRN representation obtained by our proposed method, our aim is to improve the performance of biological downstream tasks such as patient hazard prediction and disease subtype classification (graph-level task), and gene function classification (node-level task). We applied SupGCL on real GRN datasets derived from patients with multiple types of cancer, and in all experiments SupGCL achieves better performance than state-of-the-art baselines.
@@ 1 Introduction
Graph representation learning has recently attracted attention in various fields to learn a meaningful latent space to represent the connectivity and attributes in given graphs  [1]. Applications of graph representation learning are advancing in numerous areas where network data exists, such as analysis in social networks, knowledge graphs  [2, 3], and biological network analysis in bioinformatics  [4, 5].
@@
Among these applications, the use of graph representation learning for Gene Regulatory Networks (GRNs), where each node and edge represents important intracellular functions and/or processes, is particularly significant in the fields of biology and drug discovery, as it is expected to contribute to identifying therapeutic targets and understanding disease mechanisms. GRN representation learning has been applied to tasks such as inferring transcription factors  [6] and predicting drug responses in cancer cell lines  [7].
With advancements in gene expression measurement and analysis technologies, identification methods for GRNs from expression data are also evolving. Traditional GRN identification constructs networks using statistical techniques applied to patient populations. Recently, it has become possible to construct GRNs specific to individual patients, highlighting distinct gene regulatory patterns compared to the population as a whole  [8]. Hereafter in this paper, we refer to such individualized networks simply as GRNs. Similarly, in cell-based experiments such as gene knockdowns, it is now possible to estimate distinct GRNs for each experiment depending on the expression profile.
Graph representation learning applied to GRNs is believed to be effective for a wide range of biological applications. Among such methods, Graph Contrastive Learning (GCL) has gained traction. GCL enhances graph data via artificial perturbations applied to nodes or edges, and learns useful graph embeddings by maximizing the similarity between differently augmented views of the same graph  [9]. For example, we can perform contrastive learning by artificial perturbations such as randomly removing nodes to augment the GRN and consequently learn a representation for the GRN.
However, a major challenge arises in that the artificial perturbations employed in conventional GCL methods significantly deviate from genuine biological perturbations, making it difficult to learn effective representations when applied to GRNs. A related issue occurs in heterogeneous node networks, where random perturbations to nodes or edges during augmentation can disrupt topology and attribute integrity, ultimately hindering representation learning  [10, 11]. This problem is especially relevant to GRNs, where node heterogeneity exists—for example, in the presence of master regulators  [12].
@@
To address these issues, we propose a novel supervised GCL method (SupGCL) that leverages gene knockdown perturbations within GRNs. Our method uses experimental data from actual gene knockdowns as supervision, enabling biologically faithful representation learning. In gene knockdown experiments, the expression of specific genes is suppressed, representing biological perturbations that allow for the inference of GRNs. By using these perturbations as supervision signals for GCL, we can perform data augmentation that retains biological characteristics. Moreover, since our method naturally extends traditional GCL models in the direction of supervised augmentation within a probabilistic framework, conventional GCL approaches emerge as special cases of our proposed model.
@@
Our implementation and all experimental codes are available on http://github.com/xxxxxx.
2 Related Work
Graph Contrastive Learning (GCL) has inspired the development of numerous methods, largely based on the design of data augmentations and the construction of positive/negative sample pairs  [10]. GCL methods can be broadly categorized into three types according to how they generate these training pairs: (1) graph-level pairs,(2) node-level pairs, and (3) cross-model pairs.
A representative method using graph-level pairs is GraphCL [13], which applies random data augmentations to graphs and treats the resulting two graph views as a positive pair. This approach enhances the model’s ability to capture graph-level representations. However, it has been pointed out that node-level information can become obscured in the process  [14]. This limitation is especially problematic in applications like Gene Regulatory Networks (GRNs), where the semantics of individual nodes are crucial.
In contrast, methods such as GRACE, which generate node-level pairs, apply augmentations to graphs and treat embeddings of different nodes as negative pairs during training [15]. This enables more precise representation learning that captures local structure and attribute information at the node level. Although GRACE has been applied to real GRNs [6], the augmentation strategies used do not incorporate the specific biological characteristics of GRNs.
Beyond direct graph manipulations, methods like BGRL [16] generate positive pairs across models—between two instances of the same graph embedding model operating at different learning speeds—rather than relying on heuristic graph augmentations. Such cross-model pairing strategies have attracted attention for their ability to learn graph representations in a more natural manner. Notably, the recently proposed SGRL [17] achieves stable and high-performance self-supervised learning by using a pair of models: one that distributes node embeddings uniformly on a hypersphere, and another that incorporates graph topology information. However, these methods are constrained by their inability to design task-specific perturbations, and their applicability to biological networks such as GRNs remains unexplored and unverified.
@@ 3.1 Background of Graph Contrastive Learning
Although there are various definitions of contrastive learning, it can be expressed using a probabilistic model based on KL divergence over pairs of augmentations or node instances  [18]. Let $\mathcal{X}$ denote a set of entities and let $(x,y)\in\mathcal{X}\times\mathcal{X}$ be a pair from that set. The contrastive loss is formulated as follows:
@@
Here, $q_{\phi}(y|x)$ is the probability distribution of the target model with parameter $\phi$, and $p_{\theta}(y|x)$ is a reference distribution. To avoid trivial solutions when training both $p_{\theta}$ and $q_{\phi}$ simultaneously, the reference distribution $p_{\theta}$ is almost fixed. The reference model $p_{\theta}(y|x)$ is often designed as a probability that assigns a non-zero constant to positive pairs $(x,y)$ and zero to negative pairs $(x,y)$.
Graph Contrastive Learning (GCL) handles the target model $q_{\phi}(j|i)$ corresponding to a pair of nodes $(i,j)$. Consider graph operations for augmentation, order them, and represent the index of these operations by $a$. Let $z^{a}_{i}\in\mathbb{R}^{d}$ be the graph embedding of the $i$-th node obtained from the Graph Neural Network under the $a$-th augmentation operation. For two augmentation operations $(a,b)$, the pair of probability models $(p,q_{\phi}^{a,b})$ used in GCL is defined by
$p(j|i)\triangleq\delta_{ij},\quad q^{a,b}_{\phi}(j|i)\triangleq\frac{\exp(\mathrm{sim}(z_{i}^{a},z_{j}^{b})/\tau_{\rm n})}{\sum_{k\in\mathcal{V}}\exp(\mathrm{sim}(z_{i}^{a},z_{k}^{b})/\tau_{\rm n})}.$ (2)
@@
This encourages the embeddings at the node level $z_{i}^{a}$ and $z_{i}^{b}$ of the same node under different augmentation operations to be close to each other. Typically, augmentation operations $a,b$ are chosen by uniform sampling from a set of candidates $\mathcal{A}$. Hence, in practice, the expected value is minimized under the uniform distribution ${\rm U}_{\mathcal{A}}$ over $\mathcal{A}$:
@@ 3.2 Notation and Problem Definition
In this study, we describe a GRN as a directed graph $\mathcal{G}\triangleq(\mathcal{V},\mathcal{E},X^{\mathcal{V}},X^{\mathcal{E}})$ that contains information on nodes and edges. Here, $\mathcal{V}$, and $\mathcal{E}$ are the sets of nodes and edges, respectively, and each node represents a gene. $X_{i}^{\mathcal{V}}$ is the feature of the $i$-th gene, and $X_{i}^{\mathcal{E}}\in\mathbb{R}^{|\mathcal{E}|}$ is the feature of the $i$-th edge in the network. The augmentation operation corresponding to the knockdown of the $i$-th gene is modeled by setting the feature of the $i$-th gene to zero and also setting the features of all edges connected to the $i$-th gene to zero.
@@
$Z^{a}\triangleq f_{\phi}(\mathcal{G}_{a})\in\mathbb{R}^{|\mathcal{V}|\times d},\quad Y^{a}\triangleq f_{\phi}(\mathcal{H}_{a})\in\mathbb{R}^{|\mathcal{V}|\times d},$ (5)
where $z_{i}^{a}$ and $y_{i}^{a}$ denote the embedding vectors of the $i$-th node in $Z^{a}$ and $Y^{a}$, respectively, and $d$ is the embedding dimension. Note that the same GNN $f_{\phi}$ is used to produce both $Z^{a}$ and $Y^{a}$.
In this work, we train the neural network $f_{\phi}$ using the set of pairs $\{(Z^{a},Y^{a})\}_{a\in\mathcal{K}}$, where $(Z^{a},Y^{a})$ corresponds to the graph embedding obtained by the GRN augmentation operation and the embedding of the teacher GRN for the corresponding gene knockdown.
@@ 4 Method
Figure 1: Schematic overview of SupGCL (Supervised Graph Contrastive Learning). The proposed method leverages two complementary types of graph augmentations. First, it generates artificially perturbed GRNs by simulating gene knockdowns, where node and edge features are masked based on the targeted gene. Second, it incorporates biologically grounded augmentations derived from real gene knockdown experiments conducted on cancer cell lines, serving as teacher GRNs. Embeddings are extracted using a shared GNN, and both node-level and augmentation-level contrastive losses are computed via KL divergence. This biologically grounded contrastive framework enables more faithful and effective representation learning of GRNs.
For the set of embedded representations $\{(Z^{a},Y^{a})\}_{a\in\mathcal{K}}$, we consider the pair of augmentation operations $(a,b)$ and the pair of nodes $(i,j)$ according to the contrastive learning scheme. First, for the pair of augmentation operations $(a,b)$, we clarify the supervised learning problem for augmentation operations using KL divergence and then propose SupGCL using a distribution over pairs of combinations of nodes and extension operations. A sketch of the proposed method is shown in Figure 1.
@@
$p_{\phi}(b|a)\triangleq\frac{\exp\left(\mathrm{sim}_{F}(Y^{a},Y^{b})/\tau_{\rm a}\right)}{\sum_{c\in\mathcal{K}}\exp\left(\mathrm{sim}_{F}(Y^{a},Y^{c})/\tau_{\rm a}\right)},\quad q_{\phi}(b|a)\triangleq\frac{\exp\left(\mathrm{sim}_{F}(Z^{a},Z^{b})/\tau_{\rm a}\right)}{\sum_{c\in\mathcal{K}}\exp\left(\mathrm{sim}_{F}(Z^{a},Z^{c})/\tau_{\rm a}\right)},$ (6)
where $\mathrm{sim}_{F}(\cdot,\cdot)$ denotes the Frobenius inner product, and $\tau_{\rm a}>0$ is a temperature parameter. Unlike node-level learning, $p_{\phi}(b|a)$ is not a fixed constant but rather a reference distribution based on the supervised embeddings $\{Y^{a}\}_{a\in\mathcal{K}}$. Both probability models $p_{\phi}$ and $q_{\phi}$ are parameterized by the same GNN $f_{\phi}$.
Using these probability distributions, substituting the reference model $p_{\phi}(b|a)$ and the target model $q_{\phi}(b|a)$ into the formulation of contrastive learning in (1) yields the loss function for augmentation operations:
@@
To address this issue, here we first introduce a reference model $p_{\phi}(j,b|i,a)$ and a target model $q_{\phi}(j,b|i,a)$ that use conditional probabilities for each pair of node and augmentation $(i,a),(j,b)\in\mathcal{V}\times\mathcal{K}$. By substituting these into the contrastive learning formulation in (1), we derive the loss function of Supervised Graph Contrastive Learning:
@@
Furthermore, the following theorem shows that by assuming independence between nodes and augmentation operations in the reference distribution $p_{\phi}$, we can avoid the trivial solution:
@@
$\mathrm{Loss}_{\rm SupGCL}=\mathrm{E}_{a,b\sim p_{\phi}(b|a)\mathrm{U}_{\mathcal{K}}(a)}\big{[}{\rm Loss}_{\rm node}^{a,b}\big{]}+\mathrm{Loss}_{\rm Aug}.$ (9)
@@
The first term in Theorem 1 corresponds to the expectation of the node-level GCL loss $\mathrm{Loss}_{\rm node}^{a,b}$ (as defined in Equation 3) with respect to the supervised augmentation distribution $p_{\phi}(b|a)$. This allows node-level contrastive learning to reflect biological similarity between knockdown operations. Importantly, since the theorem is independent of the specific choice of the node-level model $(p,q_{\phi}^{a,b})$, any contrastive loss described by KL divergence can be used in practice. Meanwhile, the second term reduces the distributional difference between the artificially generated augmentation-based GRN and the teacher GRN. Together, these two components ensure both expressive node representations and biologically meaningful augmentations.
Moreover, the performance of node-level representation learning and the biological validity following the teacher data for augmentation operations can be controlled by the temperature parameters $\tau_{\rm n},\tau_{\rm a}$ of each probability model. In particular, when the temperature parameter $\tau_{\rm a}$ involved in the augmentation operation is sufficiently large, the augmentation operation becomes independent of the teacher GRNs $\{Y_{a}\}_{a\in\mathcal{K}}$, and coincides with the conventional node-level GCL loss function.
@@
As $\tau_{\rm a}\rightarrow\infty$, we have $p_{\phi}(b|a)\rightarrow\mathrm{U}_{\mathcal{K}}(b)$ and $q_{\phi}(b|a)\rightarrow\mathrm{U}_{\mathcal{K}}(b)$. Therefore, the expectation term becomes: $\lim_{\tau_{\rm a}\rightarrow\infty}\mathrm{E}_{a,b\sim p_{\phi}(b|a)\mathrm{U}_{\mathcal{K}}(a)}\big{[}{\rm Loss}_{\rm node}^{a,b}\big{]}=\mathrm{E}_{a,b\sim{\rm U}_{\mathcal{K}}}[\mathrm{Loss}_{\rm node}^{a,b}]$, $\lim_{\tau_{\rm a}\rightarrow\infty}D_{\mathrm{KL}}(p_{\phi}(b|a)|q_{\phi}(b|a))=0$ thus proving the corollary. ∎

### v4

@@ Abstract
Graph Contrastive Learning (GCL) is a powerful self-supervised learning framework that performs data augmentation through graph perturbations, with growing applications in the analysis of biological networks such as Gene Regulatory Networks (GRNs). The artificial perturbations commonly used in GCL, such as node dropping, induce structural changes that can diverge from biological reality. This concern has contributed to a broader trend in graph representation learning toward augmentation-free methods, which view such structural changes as problematic and to be avoided. However, this trend overlooks the fundamental insight that structural changes from biologically meaningful perturbations are not a problem to be avoided but a rich source of information, thereby ignoring the valuable opportunity to leverage data from real biological experiments. Motivated by this insight, we propose SupGCL (Supervised Graph Contrastive Learning), a new GCL method for GRNs that directly incorporates biological perturbations from gene knockdown experiments as supervision. SupGCL is a probabilistic formulation that continuously generalizes conventional GCL, linking artificial augmentations with real perturbations measured in knockdown experiments and using the latter as explicit supervisory signals. To assess effectiveness, we train GRN representations with SupGCL and evaluate their performance on downstream tasks. The evaluation includes both node-level tasks, such as gene function classification, and graph-level tasks on patient-specific GRNs, such as patient survival hazard prediction. Across 13 tasks built from GRN datasets derived from patients with three cancer types, SupGCL consistently outperforms state‑of‑the‑art baselines.
@@ 1 Introduction
Graph representation learning has recently attracted attention in various fields to learn a meaningful latent space to represent the connectivity and attributes in given graphs  (Ju et al., 2024). Applications of graph representation learning are advancing in numerous areas where network data exists, such as analysis in social networks, knowledge graphs (Hu et al., 2023; Shen & Zhang, 2023), and biological network analysis in bioinformatics  (Liu et al., 2023; Wu et al., 2021).
@@
The application of graph representation learning to Gene Regulatory Networks (GRNs), which contain information about intracellular functions and processes, is particularly important in the fields of biology and drug discovery. It is expected to contribute to the identification of therapeutic targets and the elucidation of disease mechanisms. Representation learning for GRNs has been applied to tasks such as transcription factor inference (Yu et al., 2025) and predicting drug responses in cancer cell lines (Liu et al., 2022). Advances in gene expression measurement and analysis technologies have enabled the construction of patient-specific GRNs, highlighting gene regulation patterns that differ from the population as a whole (Nakazawa et al., 2021). Hereafter, this paper will refer to such individualized networks simply as GRNs.
Among the various graph representation learning methods, Graph Contrastive Learning (GCL) has emerged as a powerful self-supervised learning framework (You et al., 2020). GCL learns by maximizing the similarity between node representations across different views of the same graph generated via data augmentation, a process that assumes the preservation of essential features like graph topology (Zhu et al., 2021; Wang et al., 2024). While augmentation methods have been refined in the application of GCL to GRNs, there have been concerns that conventional artificial perturbations like node dropping can cause structural changes so significant—disrupting the function of networks including critical nodes like master regulators—that they hinder learning (Paull et al., 2021). In response to these challenges, Augmentation-Free approaches have been developed, which perturb model parameters instead of the graph structure itself, and have shown high performance (Thakoor et al., 2021; He et al., 2024).
However, this trend overlooks a fundamental insight: that the representational shifts caused by graph augmentation are not an obstacle to overcome, but rather a rich source of information to be exploited. Consequently, it ignores the valuable opportunity to leverage data obtained from actual biological experiments, such as knockdown experiments.
@@
To address these issues, we propose a novel supervised GCL method (SupGCL) that leverages gene knockdown perturbations within GRNs. Our method uses experimental data from actual gene knockdowns as supervision, enabling biologically faithful representation learning. In gene knockdown experiments, the suppression of specific genes causes biological perturbations, leading to the observation of a new, altered GRN. By using these perturbations as supervision signals for GCL, we can perform data augmentation that retains biological characteristics. Moreover, since our method naturally extends traditional GCL models in the direction of supervised augmentation within a probabilistic framework, conventional GCL approaches emerge as special cases of our proposed model.
@@
2 Related Works
2.1 Augmentation-Free Graph Contrastive Learning (GCL)
GCL is a powerful framework for learning representations by maximizing the similarity between multiple ”views” of a graph generated through data augmentation. While early node-level methods like GRACE achieved high performance (Zhu et al., 2020), graph-level approaches like GraphCL struggled with the loss of key node features (You et al., 2021; Sun et al., 2025).
A common challenge for these methods was their heavy reliance on the choice of data augmentations. This led to a trend in augmentation-free GCL, pioneered by BGRL (Thakoor et al., 2021), which used a bootstrapping mechanism to circumvent heuristic augmentations that could damage the graph’s structure, influencing subsequent research such as simGRACE (Xia et al., 2022) and AFGRL (Lee et al., 2022). SGRL (He et al., 2024) further advanced this trend by integrating concepts like feature uniformity from early GCL to prevent the representation collapse problem often faced by methods like BGRL, achieving stable and high-performance self-supervised learning. However, these methods view perturbations to the graph as a problem to be avoided, thereby overlooking the opportunity to utilize real-world supervisory signals that arise from structural changes.
2.2 Representation Learning for Gene Regulatory Networks(GRNs)
Applying Graph Neural Networks (GNNs) to known GRNs is a growing area of research for downstream tasks such as gene function classification and patient survival prediction(Zohari & Chehreghani, 2025). The prevailing approach in this field has been supervised learning without pre-training (Liu et al., 2022; Yu et al., 2025). While link prediction addresses the important task of inferring the unknown structure of GRNs (Huynh-Thu et al., 2010; Yu et al., 2025), our work focuses on learning representations from known GRN structures for downstream applications.
Recently, self-supervised methods have been introduced to this domain. Initial efforts focused on applying general augmentation-free methods, such as the Graph Autoencoder (GAE), which learns representations by reconstructing the graph’s structure (Jung et al., 2024). More recently, the focus has begun to shift towards specialized GCL frameworks tailored for GRNs, although such approaches are still uncommon. A notable example is MuSe-GNN (Liu et al., 2023), which leverages different data modalities (e.g., scRNA-seq and scATAC-seq) as natural ”views,” avoiding the need for artificial augmentations by using biologically diverse information. In contrast to these methods, our work, SupGCL, uniquely incorporates real-world perturbations from within a single modality as supervisory signals for GCL.
@@ 3.1 Background of Graph Contrastive Learning
Although there are various definitions of contrastive learning, it can be expressed using a probabilistic model based on KL divergence over pairs of augmentations or node instances  (Alshammari et al., 2025). Let $\mathcal{X}$ denote a set of entities and let $(x,y)\in\mathcal{X}\times\mathcal{X}$ be a pair from that set. The contrastive loss is formulated as follows:
@@
Here, $q_{\phi}(y|x)$ is the probability distribution of the learned model with parameter $\phi$, and $p_{\theta}(y|x)$ is a reference distribution. To avoid trivial solutions when training both $p_{\theta}$ and $q_{\phi}$ simultaneously, the reference distribution $p_{\theta}$ is almost fixed. The reference model $p_{\theta}(y|x)$ is often designed as a probability that assigns a non-zero constant to positive pairs $(x,y)$ and zero to negative pairs $(x,y)$.
Graph Contrastive Learning (GCL) handles the learned model $q_{\phi}(j|i)$ corresponding to a pair of nodes $(i,j)$. Consider graph operations for augmentation, order them, and represent the index of these operations by $a$. Let ${\bm{z}}^{a}_{i}\in\mathbb{R}^{d}$ be the graph embedding of the $i$-th node obtained from the Graph Neural Network under the $a$-th augmentation operation. For two augmentation operations $(a,b)$, the pair of probability models $(p,q_{\phi}^{a,b})$ used in GCL is defined by
$p(j|i)\triangleq\delta_{ij},\quad q^{a,b}_{\phi}(j|i)\triangleq\frac{\exp(\mathrm{sim}({\bm{z}}_{i}^{a},{\bm{z}}_{j}^{b})/\tau_{\rm n})}{\sum_{k\in\mathcal{V}}\exp(\mathrm{sim}({\bm{z}}_{i}^{a},{\bm{z}}_{k}^{b})/\tau_{\rm n})}.$ (2)
@@
This encourages the embeddings at the node level ${\bm{z}}_{i}^{a}$ and ${\bm{z}}_{i}^{b}$ of the same node under different augmentation operations to be close to each other. Typically, augmentation operations $a,b$ are chosen by uniform sampling from a set of candidates $\mathcal{A}$. Hence, in practice, the expected value is minimized under the uniform distribution ${\rm U}_{\mathcal{A}}$ over $\mathcal{A}$:
@@ 3.2 Notation and Problem Definition
In this study, we describe a GRN as a directed graph $\mathcal{G}\triangleq(\mathcal{V},\mathcal{E},{\bm{X}}^{\mathcal{V}},{\bm{X}}^{\mathcal{E}})$ that contains information on nodes and edges. Here, $\mathcal{V}$, and $\mathcal{E}$ are the sets of nodes and edges, respectively, and each node represents a gene. ${\bm{X}}_{:,i}^{\mathcal{V}}$ is the feature of the $i$-th gene, and ${\bm{X}}_{:,i}^{\mathcal{E}}$ is the feature of the $i$-th edge in the network. The augmentation operation corresponding to the knockdown of the $i$-th gene is modeled by setting the feature of the $i$-th gene to zero and also setting the features of all edges connected to the $i$-th gene to zero.
@@
${\bm{Z}}^{a}\triangleq f_{\phi}(\mathcal{G}_{a})\in\mathbb{R}^{|\mathcal{V}|\times d},\quad{\bm{Y}}^{a}\triangleq f_{\phi}(\mathcal{H}_{a})\in\mathbb{R}^{|\mathcal{V}|\times d},$ (5)
where ${\bm{z}}_{i}^{a}$ and ${\bm{y}}_{i}^{a}$ denote the embedding vectors of the $i$-th node in ${\bm{Z}}^{a}$ and ${\bm{Y}}^{a}$, respectively, and $d$ is the embedding dimension. Note that the same GNN $f_{\phi}$ is used to produce both ${\bm{Z}}^{a}$ and ${\bm{Y}}^{a}$.
In this work, we train the neural network $f_{\phi}$ using the set of pairs $\{({\bm{Z}}^{a},{\bm{Y}}^{a})\}_{a\in\mathcal{K}}$, where $({\bm{Z}}^{a},{\bm{Y}}^{a})$ corresponds to the graph embedding obtained by the GRN augmentation operation and the embedding of the teacher GRN for the corresponding gene knockdown.
Figure 1: Schematic overview of SupGCL. Artificial augmentations are generated by simulating gene knockdowns in a patient GRN, while the teacher GRNs for supervision are derived from real-world knockdown experiments. Embeddings are extracted using a shared GNN, and both node-level and augmentation-level contrastive losses are computed via KL divergence.
@@ 4 Method
For the set of embedded representations $\{({\bm{Z}}^{a},{\bm{Y}}^{a})\}_{a\in\mathcal{K}}$, we consider the pair of augmentation operations $(a,b)$ and the pair of nodes $(i,j)$ according to the contractive learning scheme. First, for the pair of augmentation operations $(a,b)$, we clarify the supervised learning problem for augmentation operations using KL divergence and then propose SupGCL using a distribution over pairs of combinations of nodes and extension operations. A sketch of the proposed method is shown in Figure 1.
@@
$p_{\phi}(b|a)\triangleq\frac{\exp\left(\mathrm{sim}_{F}({\bm{Y}}^{a},{\bm{Y}}^{b})/\tau_{\rm a}\right)}{\sum_{c\in\mathcal{K}}\exp\left(\mathrm{sim}_{F}({\bm{Y}}^{a},{\bm{Y}}^{c})/\tau_{\rm a}\right)},\quad q_{\phi}(b|a)\triangleq\frac{\exp\left(\mathrm{sim}_{F}({\bm{Z}}^{a},{\bm{Z}}^{b})/\tau_{\rm a}\right)}{\sum_{c\in\mathcal{K}}\exp\left(\mathrm{sim}_{F}({\bm{Z}}^{a},{\bm{Z}}^{c})/\tau_{\rm a}\right)},$ (6)
where $\mathrm{sim}_{F}(\cdot,\cdot)$ denotes the cosine similarity via the Frobenius inner product, and $\tau_{\rm a}>0$ is a temperature parameter. Unlike node-level learning, $p_{\phi}(b|a)$ is not a fixed constant but rather a reference distribution based on the supervised embeddings $\{{\bm{Y}}^{a}\}_{a\in\mathcal{K}}$. Both probability models $p_{\phi}$ and $q_{\phi}$ are parameterized by the same GNN $f_{\phi}$.
Using these probability distributions, substituting the reference model $p_{\phi}(b|a)$ and the learned model $q_{\phi}(b|a)$ into the formulation of contrastive learning in (1) yields the loss function for augmentation operations:
@@
To address this issue, here we first introduce a reference model $p_{\phi}(j,b|i,a)$ and a learned model $q_{\phi}(j,b|i,a)$ that use conditional probabilities for each pair of node and augmentation $(i,a),(j,b)\in\mathcal{V}\times\mathcal{K}$. By substituting these into the contrastive learning formulation in (1), we derive the loss function of Supervised Graph Contrastive Learning:
@@
Since data augmentation affects the entire graph, we assume that the graph-level teacher distribution $p_{\phi}(b|a)$, which evaluates the similarity between augmentation operations, and the node-level teacher distribution $p(j|i)$, which evaluates node identity, are independent. This leads to the following theorem.
@@
$\mathrm{Loss}_{\rm SupGCL}=\mathbb{E}_{a,b\sim p_{\phi}(b|a)\mathrm{U}_{\mathcal{K}}(a)}\big[{\rm Loss}_{\rm node}^{a,b}\big]+\mathrm{Loss}_{\rm Aug}.$ (9)
@@
The first term in Theorem 1 corresponds to the expectation of the node-level GCL loss $\mathrm{Loss}_{\rm node}^{a,b}$ (as defined in (3)) with respect to the supervised augmentation distribution $p_{\phi}(b|a)$. This allows node-level contrastive learning to reflect biological similarity between knockdown operations. Importantly, since the theorem is independent of the specific choice of the node-level model $(p,q_{\phi}^{a,b})$, any contrastive loss described by KL divergence can be used in practice. Meanwhile, the second term reduces the distributional difference between the artificially generated augmentation-based GRN and the teacher GRN. Together, these two components ensure both expressive node representations and biologically meaningful augmentations.
Moreover, the performance of node-level representation learning and the biological validity following the teacher data for augmentation operations can be controlled by the temperature parameters $\tau_{\rm n}$ and $\tau_{\rm a}$ of each probability model. In particular, when the temperature parameter $\tau_{\rm a}$ involved in the augmentation operation is sufficiently large, the augmentation operation becomes independent of the teacher GRNs $\{{\bm{Y}}_{a}\}_{a\in\mathcal{K}}$, and coincides with the conventional node-level GCL loss function.
@@
As $\tau_{\rm a}\rightarrow\infty$, we have $p_{\phi}(b|a)\rightarrow\mathrm{U}_{\mathcal{K}}(b)$ and $q_{\phi}(b|a)\rightarrow\mathrm{U}_{\mathcal{K}}(b)$. Therefore, the expectation term becomes: $\lim_{\tau_{\rm a}\rightarrow\infty}\mathbb{E}_{a,b\sim p_{\phi}(b|a)\mathrm{U}_{\mathcal{K}}(a)}\big[{\rm Loss}_{\rm node}^{a,b}\big]=\mathbb{E}_{a,b\sim{\rm U}_{\mathcal{K}}}[\mathrm{Loss}_{\rm node}^{a,b}]$, $\lim_{\tau_{\rm a}\rightarrow\infty}D_{\mathrm{KL}}(p_{\phi}(b|a)|q_{\phi}(b|a))=0$ thus proving the corollary. ∎